# Planning In Natural Language Improves LLM Search For Code Generation

### v1

@@ 5.1 Datasets
We evaluate our search methods on three benchmarks: MBPP+, HumanEval+  [26], and LiveCodeBench [22]. MBPP [3] and HumanEval [13] are some of the most widely used code benchmarks in the field. However, since both benchmarks provide only a few test cases, [26] updates both benchmarks with additional test cases that increase the benchmarks’ robustness to reward hacking. LiveCodeBench is a benchmark for coding that consists of competitive programming problems which typically require advanced reasoning capabilities. Given the reality that coding data is often highly upsampled during pre-training [30, 15], LiveCodeBench differentiates itself from other benchmarks by taking care to segregate problems by date to avoid data contamination concerns. For this paper, we use only the subset of problems between May 2024 and September 2024 to avoid possibilities of contamination. We choose May 2024 as the cutoff date to ensure that our results with our best performing model (Claude 3.5 Sonnet) are not due to contamination, because Claude 3.5 Sonnet has a knowledge cutoff of April 2024. To ensure fair comparison, we use the same cutoff for all models evaluated, even though the precise cutoff dates for other models may vary slightly from May 2024.
@@ 5.2 Experiment Details
All models are run with temperature $0.9$ and top-$p$ of $0.95$. Temperature was determined through a coarse hyperparameter sweep on Repeated Sampling and IdeaSearch from $T\in\{0.0,0.1,0.2,\dots,1.2\}$, which we describe in Appendix F.
Both Repeated Sampling and IdeaSearch generate exactly $n$ codes, whereas PlanSearch generates a variable number of codes, usually ranging on the order of 300 to 400. To compute pass@k, we use the unbiased estimator in Equation 4 [13] [^1] Note that the estimator in Equation 4 theoretically requires that the number of successes follows a binomial distribution. Repeated Sampling and IdeaSearch obey this, but PlanSearch generations may not be independent. See Appendix N for more discussion.. If $k>n$, we assume the remaining generations did not pass. To compute pass@k for filtering, we limit the pool of codes to those that are filtered, meaning that both $n$ and $c$ may shrink in size. This can be thought of as a conditional probability, where the condition is that the code passes public tests.
@@ 5.3 Results
Our summarized results for Repeated Sampling, IdeaSearch, and PlanSearch can be found in Table 1, Figure 1, and Figure 5. Additionally, we plot our full pass@k curves for all methods, models, and datasets in Appendix A. For sake of easy comparison, we also plot all relative gains compared to Repeated Sampling@1 averaged over all models in Appendix C. For a compute-normalized comparison between Repeated Sampling and PlanSearch, see Figure 18.
@@ 6 Analysis
Our results suggest that both PlanSearch and IdeaSearch outperform basic sampling by a wide margin (Figures 121314), with PlanSearch achieving the best score across all methods and models considered. We show the detailed pass@k results for each dataset in Figures 78 and 9. We also compare with Chain-of-Thought [41] in Appendix E. Interestingly, we find that IdeaSearch performs somewhat better, which we speculate comes from differences in splitting solution sketch into two model responses, instead of doing both chain-of-thought and code solution in one model response.
Investigating the differences in specific models, we notice that trends exhibited by the pass@k curves are not uniform across all models; in fact, each curve seems unique. We hypothesize that these differences are in part due to changes in idea diversity, as investigated in Figures 62627. From the figures, we can see that our approximate diversity score accounts for much of the variance we see in the relative improvement that arrives from scaling-up inference-time compute. This correlation holds across all methods and models on the same dataset, thus suggesting that diversity score can be used as a proxy to predict for relative pass@k improvement. For further discussion on the specifics of the diversity score, see Section 6.1.
@@ 6.1 Measuring Diversity
Figure 6: We observe a strong positive correlation between the measured amount of idea diversity in a search algorithm and the resulting improvements due to search (Section 6.1). Diversity score is the probability that GPT-4o-mini believes two randomly selected output codes implement different ideas (higher is more diverse). Our findings suggest that diversity in idea space is essential for effective LLM search.
We find that diversity as measured in idea space is highly predictive of search performance, as measured by the relative improvement between a model/method’s pass@1 and its pass@200 (Figure 6). While the most common measure of diversity is entropy [36], entropy is insufficient for a number of reasons for the precise setting of LLMs [21, 45]. As a simple example, consider two different language models, one of which generates minor variations of the same program while another generates a variety of programs with different underlying ideas. Even if both models have the same entropy, the latter model will be significantly better when augmented with search capabilities.
In our setting, we measure diversity by grounding it in idea space using a simple pair-matching strategy across all generated programs. Formally, suppose we have a pool of $n$ code generations, $\{c_{1},\dots,c_{n}\}$. We assume that each piece of code implements some sketch, which can be thought to exist in some latent ‘idea’ space. We consider two sketches similar if they are within $\epsilon$ of each other in this latent space, for some choice of $\epsilon$. As such, in this space, $c_{i}$ having a similar idea to $c_{j}$ and similarly for $c_{j}$ and $c_{k}$ does not imply $c_{i}$ and $c_{k}$ share a similar idea.
To compute the diversity of such a given generation pool, we ask an LLM to judge the similarity of two ideas in the following manner. First, we construct each of the $\binom{n}{2}$ pairs. For each pair $(c_{i},c_{j})$, we judge (using an LLM) whether both $c_{i}$ and $c_{j}$ implement the same idea. We define this as the function $S(c_{i},c_{j})\in\{0,1\}$, which evaluates to $1$ if $c_{i}$ and $c_{j}$ implement the same idea and 00 otherwise. Our overall diversity score for a particular problem is then defined as:
@@
Models that output programs that all implement the same idea will have a score of $D=0$, while models that output completely unique programs will have a score of $D=1$. Overall, a score of $D$ implies that if two codes are chosen at random, the probability that they are the same idea (as measured by the LLM) is $D$. In Appendix M, we describe this measure in additional mathematical depth.
For a particular method, our reported diversity score is simply the diversity score over all problems in the considered dataset. For computational feasibility, for large $n$, we instead sample a subset of $40$ codes and test all pairs from that subset instead. In order to test code samples, we first backtranslate using an LLM to express the code in natural language before comparing each pair using both the code and the backtranslated idea. We detail the prompts used in Appendix J.5 and use OpenAI’s GPT-4o-mini as the supporting LLM.
@@ 7 Limitations and Future Work
While PlanSearch substantially improves diversity over idea space at inference-time, fundamentally, improvements in diversity should come at the post-training stage. This likely requires re-imagining the post-training pipeline for LLMs around search, instead of the current paradigm optimized for a single correct response. This may require both collecting high quality post-training data that is also sufficiently diverse, and new learning objectives that do not aim solely to maximize the expected reward of a given response. We are optimistic around future work to design significantly improved post-training objectives that maximize both quality and diversity and which specifically optimized to use inference-time compute to maximum effectiveness.
In terms of methodological improvements to PlanSearch, PlanSearch currently searches all leaf nodes in the search tree uniformly. Because of this, it becomes quickly intractable to go further than a couple levels deep, and in our experiments, we are only able to go two levels down the tree. Several approaches based on Monte-Carlo Tree Search (MCTS), such as Tree of Thought [43] or Reasoning as Planning [19], have suggested that some form of dynamic pruning and expansion of nodes can be very helpful. We are optimistic that PlanSearch can be further improved by such methods. Furthermore, PlanSearch is a fairly elementary method taking advantage of the paradigm that searching over a conceptual or idea space is an effective method to improve diversity, and thus, downstream task performance. It is completely feasible to search at an even higher level of abstraction than observations, which may be used to inject even more diversity into the final generated outputs.
@@
Fortunately, many filtering algorithms exist, which implicitly bring pass@k (for high $k$) to pass@1 (or lower $k$), i.e. shifting the original pass@k curve leftward. A simple example of this is public test filtering. As seen in Figure 5, pass@1 of filtered PlanSearch significantly improves upon pass@1 of base Repeated Sampling, which gets even better as $k$ increases. Moreover, most to almost all base models with public test filtering outperform their instruct model variants at pass@1, no matter the dataset (see Appendix I), where clearly base models are known to be worse, yet trading off for somewhat higher diversity. Thus, we argue that there exists a potential for a new paradigm—developing search algorithms which tradeoff pass@1 performance for much stronger pass@k performance, then filtering the promising generated solutions to extract the pass@k back into the pass@1.
Additionally, we focus on code generation in this paper and do not consider the applicability of PlanSearch to a broader set of domains. One point of importance is that the pass@k metric heavily used throughout code generation may not be as applicable to other domains and a larger focus on selecting the correct solution out of all possible generated candidates may be required, instead of merely generating the correct solution. However, with good filtering methods, which we demonstrate can be simple in nature, pass@k, for medium $k$, can be effectively brought down to pass@1, emphasizing a similar paradigm of increasing diversity, then strengthening existing filtering methods.
Finally, a natural extension of this work is training the underlying model itself on successful plans and code solutions obtained from PlanSearch. This has the potential to distill the pass@k into the pass@1—without inference-time methods like filtering—by reducing the likelihood of the model going down branches of the search tree which do not lead to correct solutions. We believe that such training is likely to significantly improve the model and look forward to future work in this direction.
@@ 8 Acknowledgements
We would like to thank Jason Wei, Miles Turpin, Sail Wang, Horace He, Kenneth Li, Celia Chen, Rahul Chalamala, Alan Wu, and Kevin Chang for their helpful comments, suggestions and discussion over the course of this project.
@@ Appendix C Average Relative Improvements
Figure 15: Average performance gain over all models of methods with public test filtering compared to Repeated Sampling@1, plotted over $k\in\{1,\dots,20\}$. Note that dotted lines are provided for reference of the base method pass@k (before filtering).
Figure 16: Average performance gain over all models of methods with public test filtering compared to Repeated Sampling@1, plotted over $k\in\{1,\dots,20\}$. Note that dotted lines are provided for reference of the base method pass@k (before filtering).
Figure 17: Average performance gain over all models of methods with public test filtering compared to Repeated Sampling@1, plotted over $k\in\{1,\dots,20\}$. Note that dotted lines are provided for reference of the base method pass@k (before filtering).
@@ Appendix G Diversity Score vs Search Improvement Plots for MBPP+ and HumanEval+
Appendix H Base Models vs. Instruct Models for Large Samples
@@
See Figures 293031 for Llama-3.1-8b pass@k comparisons.
See Figures 323334 for Llama-3.1-70b pass@k comparisons.
See Figures 33536 for DeepSeek-Coder-V2-Lite pass@k comparisons.
We also ran Llama-3.1-8b and DeepSeek-Coder-V2-Lite pass@k comparisons for $k$ up to $10,000$; see Figures 3728.
@@
Figure 29: Pass@k curves comparing Llama-3.1-8B’s base and instruct versions on MBPP+.
Figure 30: Pass@k curves comparing Llama-3.1-8B’s base and instruct versions on HumanEval+.
Figure 31: Pass@k curves comparing Llama-3.1-8B’s base and instruct versions on LiveCodeBench.
Figure 32: Pass@k curves comparing Llama-3.1-70B’s base and instruct versions on MBPP+.
Figure 33: Pass@k curves comparing Llama-3.1-70B’s base and instruct versions on HumanEval+.
Figure 34: Pass@k curves comparing Llama-3.1-70B’s base and instruct versions on LiveCodeBench.
Figure 35: Pass@k curves comparing DeepSeek-Coder-V2-Lite’s base and instruct versions on HumanEval+.
Figure 36: Pass@k curves comparing DeepSeek-Coder-V2-Lite’s base and instruct versions on LiveCodeBench.
@@
Appendix I Base Models vs. Instruct Models with Public Test Filtering
We repeat the graphs from Appendix H, but with public test filtering. We find that base models with public test filtering almost always exceed the pass@1 of their instruct model variants.
See Figures 383940 for Llama-3.1-8b pass@k comparisons with public test filtering.
See Figures 414243 for Llama-3.1-70b pass@k comparisons with public test filtering.
See Figures 444546 for DeepSeek-Coder-V2-Lite pass@k comparisons with public test filtering.
Figure 38: Pass@k curves comparing Llama-3.1-8B’s base and instruct versions on MBPP+ with public test filtering.
Figure 39: Pass@k curves comparing Llama-3.1-8B’s base and instruct versions on HumanEval+ with public test filtering.
Figure 40: Pass@k curves comparing Llama-3.1-8B’s base and instruct versions on LiveCodeBench with public test filtering.
Figure 41: Pass@k curves comparing Llama-3.1-70B’s base and instruct versions on MBPP+ with public test filtering.
Figure 42: Pass@k curves comparing Llama-3.1-70B’s base and instruct versions on HumanEval+ with public test filtering.
Figure 43: Pass@k curves comparing Llama-3.1-70B’s base and instruct versions on LiveCodeBench with public test filtering.
Figure 44: Pass@k curves comparing DeepSeek-Coder-V2-Lite’s base and instruct versions on MBPP+ with public test filtering.
Figure 45: Pass@k curves comparing DeepSeek-Coder-V2-Lite’s base and instruct versions on HumanEval+ with public test filtering.
Figure 46: Pass@k curves comparing DeepSeek-Coder-V2-Lite’s base and instruct versions on LiveCodeBench with public test filtering.
Appendix J Prompts
J.1 Backtranslation
J.1.1 Backtranslate System Prompt
@@
J.1.2 Implement Backtranslation Idea
@@
J.4 PlanSearch
J.4.1 Prompt for Observation Part 1
@@
J.4.2 Prompt for Observation Part 2
@@
J.5 Measuring Diversity
Appendix K Competitive Programming
Appendix L A Model of Repeated Sampling: Pass@k
@@
Appendix M Mathematics of the Diversity Measure
While our choice of a diversity metric is intuitive, one should note that there are a number of intriguing details that result from our definition. In particular, it is not necessarily the case that a model that outputs $k$ unique ideas out of $n$ samples to achieve a diversity score of $\frac{k}{n}$. Consider an example of $n=9$ codes, separated into $3$ cliques of $3$, where each clique implements the same idea (and separate cliques implement separate ideas). In this setup, $\frac{1}{3}$ of ideas are unique, but in our metric, there are 3 matching idea pairs (and 9 total matching idea pairs) out of (${9\choose 2}=36$, for a diversity score of $1-\frac{9}{36}=\frac{3}{4}$.
Appendix N Biased Estimator for Pass@K Due to Non-Independence of PlanSearch

### v2

@@ 5.1 Datasets
We evaluate our search methods on three benchmarks: MBPP+, HumanEval+  [28], and LiveCodeBench [24]. MBPP [3] and HumanEval [14] are some of the most widely used code benchmarks in the field. However, since both benchmarks provide only a few test cases, [28] updates both benchmarks with additional test cases that increase the benchmarks’ robustness to reward hacking. LiveCodeBench is a benchmark for coding that consists of competitive programming problems which typically require advanced reasoning capabilities.
Given the reality that coding data is often highly upsampled during pre-training [33, 16], LiveCodeBench differentiates itself from other benchmarks by taking care to segregate problems by date to avoid data contamination concerns. For this paper, we use only the subset of problems between May 2024 and September 2024 to avoid possibilities of contamination. We choose May 2024 as the cutoff date to ensure that our results with our best performing model (Claude 3.5 Sonnet) are not due to contamination, because Claude 3.5 Sonnet has a knowledge cutoff of April 2024. To ensure fair comparison, we use the same cutoff for all models evaluated, even though the precise cutoff dates for other models may vary slightly from May 2024.
@@ 5.2 Experiment Details
All models are run with temperature $0.9$ and top-$p$ of $0.95$. (o1-mini was run with temperature $1.0$ and top-$p$ of $1.0$ because of API constraints.) Temperature was determined through a coarse hyperparameter sweep on Repeated Sampling and IdeaSearch from $T\in\{0.0,0.1,0.2,\dots,1.2\}$, which we describe in Appendix F.
Both Repeated Sampling and IdeaSearch generate exactly $n$ codes, whereas PlanSearch generates a variable number of codes, usually ranging on the order of 300 to 400. To compute pass@k, we use the unbiased estimator in Equation 4 [14] [^1] Note that the estimator in Equation 4 theoretically requires that the number of successes follows a binomial distribution. Repeated Sampling and IdeaSearch obey this, but PlanSearch generations may not be independent. See Appendix O for more discussion.. If $k>n$, we assume the remaining generations did not pass. To compute pass@k for filtering, we limit the pool of codes to those that are filtered, meaning that both $n$ and $c$ may shrink in size. This can be thought of as a conditional probability, where the condition is that the code passes public tests. For more information on public test filtering, see Section 5.4.
@@ 5.3 Results
Our summarized results for Repeated Sampling, IdeaSearch, and PlanSearch can be found in Table 1, Figure 1, and Figure 5. We find that PlanSearch improves over existing methods for all models and benchmarks considered.
Additionally, we plot our full pass@k curves for all methods, models, and datasets in Appendix A. Due to prohibitively high costs, we only evaluate o1-mini on LiveCodeBench, as HumanEval+ and MBPP+ show strong saturation effects even with weaker models like GPT4o-mini. The pass@k curves of o1-mini and associated discussion can be found in Appendix K.
For sake of easy comparison, we also plot all relative gains compared to Repeated Sampling@1 averaged over all models in Appendix C. For a compute-normalized comparison between Repeated Sampling and PlanSearch, see Figure 18. We also ablate over the design choices made for PlanSearch in Appendix H.
@@ 6 Analysis
Our results suggest that both PlanSearch and IdeaSearch outperform basic sampling by a wide margin (Figures 121314), with PlanSearch achieving the best score across all methods and models considered. Since o1-mini is unique from all other models tested, we show and discuss its unique results in Appendix K. We show the detailed pass@k results for each dataset in Figures 78 and 9. We also compare with Chain-of-Thought [46] in Appendix E. Interestingly, we find that IdeaSearch performs somewhat better, which we speculate comes from differences in splitting solution sketch into two model responses, instead of doing both chain-of-thought and code solution in one model response.
Investigating the differences in specific models, we notice that trends exhibited by the pass@k curves are not uniform across all models; in fact, each curve seems unique. We hypothesize that these differences are in part due to changes in idea diversity, as investigated in Figures 62627. Figure 6 includes o1-mini diversities, which also follow the observed trend. From the figures, we can see that our approximate diversity score accounts for much of the variance we see in the relative improvement that arrives from scaling-up inference-time compute. This correlation holds across all methods and models on the same dataset, thus suggesting that diversity score can be used as a proxy to predict for relative pass@k improvement. For further discussion on the specifics of the diversity score, see Section 6.1.
@@ 6.1 Measuring Diversity
Figure 6: We observe a strong positive correlation between the measured amount of idea diversity (higher score is better) in a search algorithm and the resulting improvements due to search. See Section 6.1 for information regarding the diversity score.
We find that diversity as measured in idea space is highly predictive of search performance, as measured by the relative improvement between a model/method’s pass@1 and its pass@200 (Figure 6). While entropy is a common diversity measure [40], it is oftentimes insufficient for LLM settings [23, 55]. For instance, consider two different language models, one of which generates minor variations of the same program while another generates a variety of programs with different underlying ideas. Even if both models have the same entropy, the latter model will perform significantly better if augmented with search.
In our setting, we measure diversity in idea space via pairwise comparisons across all generated programs. Let $\{c_{1},\dots,c_{n}\}$ be the pool of $n$ generated codes from a model, each corresponding to some latent idea. We consider two sketches similar if they are within $\epsilon$ of each other in this latent space, for some choice of $\epsilon$. As such, in this space, $c_{i}$ having a similar idea to $c_{j}$ and similarly for $c_{j}$ and $c_{k}$ does not imply $c_{i}$ and $c_{k}$ share a similar idea.
To compute the diversity of such a model, we construct the $\binom{n}{2}$ pairs and evaluate their similarity using an LLM. We define $S(c_{i},c_{j})\in\{0,1\}$, where $S(c_{i},c_{j})=1$ iff $c_{i}$ and $c_{j}$ are similar. Our overall diversity score for a particular problem is then defined as:
@@
A model that produces programs that all implement identical ideas will have a diversity score of $D=0$, whereas one generating codes with all unique ideas have $D=1$. Notably, a method with a score of $D$ is equivalent to the probability that two randomly selected programs from the method are similar to each other, as measured by the LLM. See Appendix Q for more details.
For a particular method, our reported diversity score is simply the diversity score averaged over all problems in the considered dataset. For large $n$, we sample a subset of $40$ codes and compare all pairs from said subset. To evaluate $S$, we first backtranslate both codes to express each code in natural language before comparing each pair using both the code and its corresponding backtranslated idea. Full prompt details can be found in Appendix O.1. We use OpenAI’s GPT-4o-mini as the supporting LLM.
@@ 7 Limitations and Future Work
While PlanSearch substantially improves diversity over idea space at inference-time, fundamentally, improvements in diversity should come at the post-training stage, like with methods such as o1 [32]. This likely requires re-imagining the post-training pipeline for LLMs around search, instead of the current paradigm optimized for a single correct response. We are optimistic about future work in designing improved post-training objectives to maximize both quality and diversity, while specifically optimized to use inference-time compute to maximum effectiveness.
@@
Fortunately, many filtering algorithms exist, which mitigates this tradeoff by implicitly bringing pass@k (for high $k$) to pass@1 (or lower $k$), i.e. shifting the original pass@k curve leftward. Even the simplest filtering—public test filtering—improves PlanSearch’s pass@1 significantly above Repeated Sampling’s pass@1, which continues as $k$ increases. Moreover, most to almost all base models with public test filtering outperform their instruct model variants at pass@1, no matter the dataset (see Appendix J). Since base models’ pass@1 is known to be worse than instruct models to trade off for higher diversity, we suggest a new paradigm—developing search algorithms which tradeoff pass@1 performance for much stronger pass@k performance, then filtering the generated solutions to extract the pass@k back into pass@1.
With good filtering methods, which we demonstrate can be simple in nature, pass@k, for medium $k$, can be effectively brought down to pass@1, emphasizing a similar paradigm of increasing diversity, then strengthening existing filtering methods, even for domains outside of code generation that are out of scope of this paper.
Finally, a natural extension of this work is training the underlying model itself on successful plans and code solutions obtained from PlanSearch. This has the potential to distill the pass@k into the pass@1—without inference-time methods like filtering—by reducing the likelihood of the model going down unfavorable branches of the search tree. We believe that such training is likely to significantly improve the model and look forward to future work in this direction.
@@ 8 Acknowledgements
We would like to thank Jason Wei, Charlie Snell, Miles Turpin, Sail Wang, Horace He, Kenneth Li, Celia Chen, Rahul Chalamala, Alan Wu, Kevin Chang, and Jerry Wei for their helpful comments, suggestions and discussion over the course of this project.
@@ Appendix C Average Relative Improvements
Figure 15: Average performance gain over all models of methods with public test filtering compared to Repeated Sampling@1, plotted over $k\in\{1,\dots,20\}$.
Figure 16: Average performance gain over all models of methods with public test filtering compared to Repeated Sampling@1, plotted over $k\in\{1,\dots,20\}$.
Figure 17: Average performance gain over all models of methods with public test filtering compared to Repeated Sampling@1, plotted over $k\in\{1,\dots,20\}$.
@@ Appendix G Diversity Score vs Search Improvement Plots for MBPP+ and HumanEval+
Appendix H Ablations
We run ablations on the core parts of PlanSearch, using GPT-4o-mini on LiveCodeBench [24]. On the diverse observation generation side, we first verify our choice of $S=2$—the maximum subset size to sample from out of a given pool of observations—and also compare performance across varying the number of observation layers used (Figures 2829). On the implementation side, we compare different sections of the proposed pipeline (see Figure 2) to translate combinations of observations to code in Figure 30. We compare each method’s pass@k from $k=1$ to $k=200$.
Figure 28: We run ablations of PlanSearch with different $S$—controlling the maximum subset size from a given pool of observations. There is not a large difference between different $S$, but we find that $1$ or $2$ is slightly more optimal, although if more completions are desired, $S$ can be increased to $3$ as well.
The effect $S$, the maximum observation subset size to build upon, has on performance is not overly significant; there are small degradations as $S$ is increased to $3$, but not noticeable. We choose $S=2$ to obtain more code completions. See Figure 28.
Figure 29: We run ablations of PlanSearch with different $L$—controlling the maximum order of observation used, i.e., how many layers the observation tree will search. We find pass@k scales with respect to increasing $L$
Increasing $L$, the maximum number of layers of the observation tree, increases pass@k at large enough $k$ (above $50$). We choose $L=2$ to strike a balance between extracting a large pass@k gain while keeping compute costs reasonable. See Figure 29.
Figure 30: We run ablations of PlanSearch with different methods of translating a combination of observations to code.
From Figure 30, we see that our overall translation step adds minor pass@k gains. We deconstruct the translation step into parts: the pseudocode step, the fix step (i.e., asking the model to fix its proposed solution sketch), and creating the solution sketch at all (which includes the fix step).
  • “No solution sketch, no pseudocode” implies using a given observation combination to directly prompt for the solution code.
  • “No pseudocode” implies skipping the pseudocode step. In other words, given a solution sketch, the sketch is directly translated into code.
  • “No fix step, no pseudocode” implies the fix step is skipped, as well as the pseudocode. In order to have the same number of completions, the whole PlanSearch pipeline is run twice.
Appendix I Base Models vs. Instruct Models for Large Samples
@@ Appendix H Ablations
See Figure 31 for all base and instruct model comparisons between DeepSeek-Coder-V2-Lite, Llama-3.1-8B, and Llama-3.1-70B on all three datasets.
We also provide Llama-3.1-8b and DeepSeek-Coder-V2-Lite pass@k comparisons for $k$ up to $10,000$; see Figures 3332.
Figure 31: Pass@k curves comparing DeepSeek-Coder-V2-Lite, Llama-3.1-8B, and Llama-3.1-70B base and instruct performance.
@@
Appendix J Base Models vs. Instruct Models with Public Test Filtering
We repeat the graphs from Appendix I, but with public test filtering. We find that base models with public test filtering almost always exceed the pass@1 of their instruct model variants.
See Figure 34 for all base and instruct model comparisons between DeepSeek-Coder-V2-Lite, Llama-3.1-8B, and Llama-3.1-70B on all three datasets with public test filtering.
We also report Llama-3.1-8b and DeepSeek-Coder-V2-Lite pass@k comparisons with public test filtering for $k$ up to $10,000$; see Figures 3536.
Figure 34: Pass@k curves comparing DeepSeek-Coder-V2-Lite, Llama-3.1-8B, and Llama-3.1-70B base and instruct performance with public test filtering.
Figure 35: Pass@k curves comparing Llama-3.1-8B’s base and instruct versions on LiveCodeBench with up to $1,000$ completions and with public test filtering.
Figure 36: Pass@k curves comparing DeepSeek-Coder-V2-Lite’s base and instruct versions on LiveCodeBench with up to $1,000$ completions and with public test filtering.
Appendix K OpenAI O1
We also do a brief analysis of OpenAI’s o1 [32] models. We run PlanSearch on top of o1-mini, which is the best o1 model available through API access at the time of writing for competitive coding. Due to both severe cost constraints and dataset saturation on HumanEval+ and MBPP+, we only run o1-mini on LiveCodeBench [24].
First, we investigate its pass@k trends on LiveCodeBench. (See Figure 37.) As expected, PlanSearch trails Repeated Sampling for low $k$. However, at high $k$, PlanSearch competes with/slightly outperforms Repeated Sampling.
We believe there to be two reasons why there is not as large of an improvement compared to other non-search-based models. First, there may be diminishing returns when stacking multiple search methods on top of each other naïvely. Second, LiveCodeBench is noticeably saturated, reaching solve-rates of over $90\%$. Thus, it is much harder to notice large jumps in dataset performance, as can be seen in pass@k curves (see Figure 7) on HumanEval+ [28]. There, Repeated Sampling, IdeaSearch, and PlanSearch all perform roughly the same, with PlanSearch only slightly outperforming the others. We believe a combination of these effects also implies a greater $k$ is required to notice a large difference between Repeated Sampling and PlanSearch.
Figure 37: Pass@k performance methods applied on top of o1-mini, on LiveCodeBench, plotted over $k\in\{1,\dots,200\}$. PlanSearch slightly outperforms Repeated Sampling at high $k$; larger $k$ is not run due to cost constraints.
We notice that the OpenAI frequently refuses to answer queries from PlanSearch, likely due to efforts to minimize revealing the full chain-of-thought the model is using. To combat this, we find that filtering out the words steps, step, and quote reduces the fraction of unanswered queries from roughly $40\%$ to less than $0.5\%$. For any query that is still refused after this approach, we route the same prompt to GPT-4o. In addition, we remove the pseudocode step to reduce the amount of flagged responses.
On the diversity end, we apply the same methodology as described in Section 6.1. The corresponding plot can be found in Figure 6. Even though o1-mini is strong on its own, we find that it is not very diverse, which is consistent with our claim that diversity correlates with relative performance gain seen with increasing $k$.
Appendix L Bar Charts
See Figures 38391. These plot pass@1 and pass@200 of select methods between Repeated Sampling and PlanSearch, on datasets MBPP+, HumanEval+ [28], and LiveCodeBench [24].
Figure 38: Bar chart with Repeated Sampling@1, Repeated Sampling@200, and PlanSearch@200, on MBPP+.
Figure 39: Bar chart with Repeated Sampling@1, Repeated Sampling@200, and PlanSearch@200, on HumanEval+.
Appendix M Prompts
M.1 Backtranslation
M.1.1 Backtranslate System Prompt
@@
M.1.2 Implement Backtranslation Idea
@@
M.4 PlanSearch
M.4.1 Prompt for Observation Part 1
@@
M.4.2 Prompt for Observation Part 2
@@
Appendix N A Model of Repeated Sampling: Pass@k
@@
Appendix O Biased Estimator for Pass@K Due to Non-Independence of PlanSearch
@@
O.1 Measuring Diversity
Appendix P Competitive Programming
Appendix Q Mathematics and Examples of the Diversity Measure
While our choice of a diversity metric is intuitive, one should note that there are a number of intriguing details that result from our definition.
For example, it is the case that with $k$ unique ideas and $n$ samples of each idea, respectively (for a total of $kn$ total generated codes), we achieve a diversity score approaching $(k-1)/k$.
For a quick proof, suppose that there are $k$ cliques, each of $n$ size. Each clique represents a unique idea. We wish to capture the number of unfilled edges over the number of possible edges as our diversity score:
$\frac{\binom{k}{2}n^{2}}{\binom{kn}{2}}=\frac{(k-1)n}{kn-1}$ (5)
which converges to $\frac{k-1}{k}$ as $n$ grows large.
Our formulation seems more intuitive than other proposals, such as one that simply counts how many unique ideas $k$ lie within a pool of size $n$ to compute $k/n$ as the diversity score.
For instance, suppose there are only two unique ideas, one clique of which is of size $12n-$ 1, and the other only output once. The simple proposal would compute both the diversity score of this uneven group and that of an even group (where both ideas are output $n$ times) to be $\frac{2}{2n}=\frac{1}{n}$.
However, our score would compute the even group to be $\frac{n}{2n-1}$, and the uneven group as:
$\frac{1}{2n}\cdot 1+\frac{2n-1}{2n}\cdot\frac{1}{2n-1}=\frac{1}{n}$ (6)
Instead of counting edges, we compute the probability that two randomly selected outputs have similar idea to each other, which is another interpretation of our diversity score.
It seems clear that a case with $12n-$ 1 instances of idea 1 and $1$ instance of idea 2 is ‘less diverse’ than a case with $n$ instances of both idea 1 and idea 2. A naïve proposal may score these two as being the same diversity, whereas our score scores them as $1/n$ and roughly $1/2$, respectively.
Appendix R Additional Directions for Future Work
In terms of methodological improvements to PlanSearch, PlanSearch currently searches all leaf nodes in the search tree uniformly. Because of this, it becomes quickly intractable to go further than a couple levels deep, and in our experiments, we are only able to go two levels down the tree. Several approaches based on Monte-Carlo Tree Search (MCTS), such as Tree of Thought [48] or Reasoning as Planning [21], have suggested that some form of dynamic pruning and expansion of nodes can be very helpful. We are optimistic that PlanSearch can be further improved by such methods.
Furthermore, PlanSearch is a fairly elementary method taking advantage of the paradigm that searching over a conceptual or idea space is an effective method to improve diversity, and thus, downstream task performance. It is completely feasible to search at an even higher level of abstraction than observations, which may be used to inject even more diversity into the final generated outputs.